For p∈P3(R) given by p(x)=a0+a1x+⋯+a3x3, let s(p)=a0+a1+a2+a3 and det(p)=a0. Also, corresponding to the polynomial p∈P3(R), we define the polynomial p∗ to be p(−x). Which of the following are subspaces of P3(R) ?

1
Expert's answer
2022-02-08T16:30:03-0500

Subspace of "P_3(R)" satisfies three conditions

a) Contain zero vector

b) closed under addition

c) closed under scalar multiplication


Zero Vector

"\forall_a\in\R" such that "a=0"

"P(R)=0" is part of set


For any "P(R):P(R)+0=P(R)"

"\therefore" The set contain zero vector


Vector addition

For any two polynomials "a_1x^3" and "a_2x^3"

"a_1x_3+a_2x_3=(a_1+a_2)x^3"

"=kx^3\in\>ax^3"


So it is closed under vector addition


Scalar multiplier

Chosing arbitrary polynomial "ax^3"

"b.ax^3=bax^3\in\>ax^3"


Therefore a subspace of "P_3(R)" is closed under scalar multiplication